OpenStudy (anonymous):

Will give medal For questions 1-4 find the simplified form of each expression A. (-5)^-4 B. (-8.3)^-1 3. -(10)^-1 4. -1/p^-6

OpenStudy (anonymous):

Is b. -8.3?\

OpenStudy (anonymous):

@zepdrix can you help me out?

OpenStudy (shaik0124):

similarly-5^-4 can be written as 1/(-5)^4 =1/-5*-5*-5*-5=1/625

OpenStudy (shaik0124):

1/625 =0.0016

OpenStudy (shaik0124):

3)-10 2)-8.3 4)-1/p^-4 =-1*p^4

OpenStudy (shaik0124):

did u understand

OpenStudy (anonymous):

yeah i think so but wouldn't 3 be -1/10?